OpenStudy (anonymous):

What is the cubed root of NEGATIVE 27/125?

OpenStudy (anonymous):

-3/5

OpenStudy (anonymous):

it's cubed kumar not sqrt so no imaginary

OpenStudy (anonymous):

oh yha!! its not imaginary... sorry...

OpenStudy (anonymous):

I agree. So it is correct as -3/5???

OpenStudy (anonymous):

yes, but is the negative given in the numerator or in up and down?

OpenStudy (anonymous):

in the middle so the fraction of 27/125 is negative

OpenStudy (anonymous):

my answer was for the negative in the numirator

OpenStudy (anonymous):

even it is the numerator or in the denominator the answer will still be negative

OpenStudy (anonymous):

yes, numirator, the answer is correct

OpenStudy (anonymous):

\[\sqrt[3]{\frac{-27}{125}}=\frac{\sqrt[3]{-27}}{\sqrt[3]{125}} = \frac{-3}{5}\]

OpenStudy (anonymous):

@kumar , I'm asking is 2 negative signs or one?

OpenStudy (anonymous):

@rspencer--the negative sign is in the middle in front of 27/125 which I know that you can move it to either the top or the bottom

OpenStudy (anonymous):

Well, I moved it to the top